### v1

@@ III Summary of Participating System Approaches
Lifelog retrieval systems must interpret user inputs across multiple modalities, such as free-text queries, visual examples, and contextual filters (e.g., time, location, or activity). The strategies for interpreting these inputs vary by system. The strategies used vary significantly across systems, ranging from rule-based filtering to embedding-based semantic search and LLM-driven query interpretation. Tran et al. [77] present a general pipeline for LSC systems, illustrating how user inputs are processed, matched against indexed data, and refined through user interaction. This model reflects the common architecture seen in many LSC submissions and helps contextualise the multimodal input-handling approaches used in the challenge.
The early years of LSC from 2018 to 2021 saw a gradual progress in lifelog retrieval methods. Developments from this period are detailed in [24] and [80], which documented the advancements made in the initial and later years of this phase respectively.
@@
This section reviews the major technical advances made in LSC’22, LSC’23, and LSC’24 and provides a structured analysis. Tables III and IV summarize the systems across these editions, highlighting key improvements in UI and UX, search and retrieval techniques, and advanced features.
@@ III-A1 LSC’22: The Rise of CLIP-Based Embedding Models
LSC’22 marked a significant shift toward embedding-based retrieval, with many teams integrating contrastive language–image pretraining (CLIP) models [59] for text-to-image search, which are language-image transformers that are trained in a contrastive way to maximize text-image similarity. Systems such as MyScéal [82] and LifeSeeker 4.0 [53] adopted CLIP to improve semantic understanding, enabling more abstract and flexible queries compared to traditional keyword-based retrieval. Memento 2.0 [1] introduced a weighted ensemble approach for CLIP integration, optimizing retrieval efficiency.
User experience improvements were also a focus. MyScéal introduced a more accessible search interface, while LifeSeeker 4.0 used event clustering to enhance navigation. vitrivr-VR [71] introduced expanded their Virtual Reality (VR) search enviroment, offering three immersive interfaces for lifelog exploration. LIFEXPLORE [44] improved semantic geo-location enrichment and search result ordering to enhance retrieval accuracy. Alternative interaction methods gained traction. Voxento 3.0 [4] implemented voice-based search with speech-to-query mechanisms, while FIRST 3.0 [35] and vitrivr [31] enhanced their retrieval pipelines by incorporating external search engines for unfamiliar concept expansion.
@@ III-B1 Embedding-based Retrieval
Embedding-based retrieval approaches have transformed lifelog search by mapping both text and images into a shared high-dimensional semantic space. Cosine similarity between the search query and the images is directly used to rank the result. Unlike traditional concept-based retrieval, which relies on explicit labels (e.g., ‘beach’ or ‘car’), embeddings capture deeper contextual meanings, enabling the retrieval of relevant data even when exact keywords are missing. For instance, users can issue abstract queries like ‘a relaxing afternoon with friends’ and receive semantically related images and videos.
This technique was mentioned in the previous analysis of LSC’21 [80], and it has become the dominant approach since LSC’22. Text and image embeddings are trained using large datasets of paired image-text data, allowing models such as CLIP [59], BLIP [47], and BLIP-2 [46] to understand complex relationships. CLIP aligns text and image embeddings to maximize their agreement for correct pairs, supporting zero-shot queries. BLIP and BLIP-2 build upon this by improving the visual question-answering and captioning capabilities, making them ideal for multimodal queries. Some optimisation methods such as KNN search [81] and FAISS [41] can be used to speed up the search as in the case of LIFEXPLORE [68] and Memento [2]. Although there is some effort in fine-tuning the embedding models [76], large-scale pre-trained models are generally more robust and are often used, either directly or in a weighted ensemble [2].
@@
While many systems relied on retrieving images individually, several recent approaches extended these embeddings to incorporate temporal context by operating over structured segments of lifelog data. Rather than treating each image in isolation, systems like MyEachtra [79], Memento 4.0 [3], and MemoriEase 2.0 [85] performed hierarchical or content-based segmentation to construct event-level representations, where visual features from temporally adjacent frames were aggregated. These event embeddings served as the atomic units for retrieval or reasoning, enabling systems to interpret queries involving longer activities or complex temporal dependencies. This shift toward event-centric representations reflects a growing interest in modeling lifelog data in ways that align more closely with human memory and real-world activity flow. Such a trend extends beyond lifelogging and is also evident in the wider multimedia retrieval community, as discussed in [67].
@@ III-B2 Large Language Models, Conversational Search, and Question Answering
The adoption of LLMs in LSC’24 marked a major milestone in lifelog retrieval. LLMs have introduced powerful language understanding capabilities to lifelog retrieval systems. By pretraining on vast corpora of text, LLMs can interpret complex, natural language queries and support dialogue-based search. In lifelog search, this means users can refine their searches through follow-up questions or clarifications, making the retrieval process more interactive and personalized.
@@ III-B3 User Interface and User Experience
  • LSC’23 saw the introduction of dynamic result sequencing and faceted filtering in LIFEXPLORE [68], enhancing search result organization. MEMORIA [60] and Memento [2] focused on detailed image annotations and user-selectable embedding models, respectively.
  • LSC’24 showcased a range of UI/UX improvements, with several teams participating in the challenge with the focus of UI/UX enhancements for different levels of user expertise, such as Retrospect [73], LifeLens[37], and VitaChronicle [57]. MyEachtraX [83] brought mobile-friendly design to lifelog search to enhance accessibility. Conversational interfaces, as mentioned earlier, are gaining in popularity, allowing users to interact with lifelog systems through natural language queries. These interfaces further enhance UX by guiding users through the search process step-by-step, allowing for query refinement through dialogue rather than manual re-entry.
Additionally, immersive interfaces gained further traction, with Virtual Reality (VR) and Augmented Reality (AR) lifelog retrieval systems offering novel ways to explore lifelog data. VR allows users to ‘step into’ their lifelogs, navigating time and space through virtual reconstructions. AR overlays lifelog information onto real-world surroundings, enabling contextual re-experiencing of past events. VR systems typically use spatial metaphors such as timelines and maps to organize lifelog entries. Head-gaze tracking and hand gestures often replace traditional input methods, making the interaction more natural. Collaborative VR environments take this further by allowing multiple users to explore lifelog data together. Some notable VR/AR systems showcased at LSC’24 include vitrivr-VR [72], which provided a map-based VR interface for lifelog exploration, and CollaXRSearch [50], which introduced a shared VR workspace for collaborative lifelog analysis. Despite the potential of VR/AR interfaces, they face adoption challenges due to hardware requirements and user comfort. However, as VR/AR technologies mature, they may redefine how users interact with lifelog archives.
@@ IV Comparative Analysis of System Performance through the Years
Since the LSC has been running for several years, it is possible to compare system performance in LSC’24 with previous editions, LSC’22 and LSC’23, which shared the same dataset and tasks (with the exception of the QA task in 2022, as the form of the answer at that time was an image, whereas in later years, it was text-based). Some systems participated across multiple years and experimented with different approaches, giving us insights on the effectiveness of various strategies. This section focuses on trends in system performance across three key metrics: correctness of submissions, time to the first correct submission, and overall scores.
@@ IV-A Performance of Systems
The evaluation metrics used in the LSC are based on the time taken to submit the correct image and the number of correct and incorrect submissions. The number of participating teams increased from 9 in 2022 to 14 in 2023, reaching 35 in 2024. In 2024, some systems had multiple instances (e.g., SYSTEM1, SYSTEM2), which were treated as independent participants. This increased the total number of submissions and led to higher performance variability.
@@ IV-A1 Correctness of Submissions
For KIS tasks, the mean number of correct submissions increased from 0.78 in 2022 to 0.87 in 2023, indicating improved system strategies. However, it dropped to 0.68 in 2024, while precision decreased to 0.64, likely due to exploratory submissions from system instances with varied configurations. Although some configurations were effective, others likely introduced noise, contributing to the lower precision. QA tasks showed consistent correct submissions (0.78 in both 2023 and 2024), but precision improved slightly from 0.50 to 0.52. Ad-hoc tasks exhibited the widest variability, with some teams performing very well and others submitting many incorrect attempts. While the median correct submissions remained stable, the spread widened in 2024, suggesting that not all configurations of the same core system performed consistently.
The overall precision for KIS and QA tasks was higher than for ad-hoc tasks, which consistently recorded lower precision due to their open-ended nature. The introduction of multiple system instances in 2024 increased the diversity of approaches but also amplified the volume of incorrect submissions, impacting precision across tasks.
@@ IV-A2 Time to First Correct Submission
Figure IV-A2 shows the distribution of the time taken by teams to make their first correct submission. For KIS and QA tasks, the median times remained relatively stable across years, but the spread increased in 2024 due to system instance variability. This indicates that while some instances responded quickly, others took considerably longer. Outliers in 2024 represent configurations that struggled to deliver timely correct submissions, demonstrating the varied effectiveness of system tuning across instances.
For QA tasks, the spread of times was consistently large across the years, but a slight improvement in the median time in 2024 suggests some gains in system efficiency. The ad-hoc retrieval tasks showed the highest variability in 2024, with some teams submitting correct answers quickly and others taking substantially longer. This wide range of performance reflects the complexity of open-ended queries and the exploratory nature of multiple instances, with different configurations demonstrating varying levels of success.
@@ IV-A3 Scores of Regular Teams by Task Type and Year
[h][width=0.99]Figures/Trends/score_by_team_and_year_regularScore of regular teams—those that participated in more than one year—in LSC’22, ’23, and ’24. Figure IV-A3 presents the scores of regular teams—those who participated in at least two editions—across KIS, QA, and Ad-hoc tasks from 2022 to 2024. The heatmap highlights year-on-year trends in performance. It is also important to note that the systems are changing every year to incorporate new approaches and techniques, which can impact performance. This analysis focuses on regular teams to understand if they have adapted and improved over time.
In KIS tasks, top-performing teams like MyScéal, LifeInsight, and LifeSeeker maintained consistently high scores across multiple years. However, LifeLens and LifeGraph saw significant score declines in 2024, potentially due to system changes or difficulties adapting to new task types. MEMORIA showed a notable improvement, increasing its score from 10 in 2022 to 90 in 2023, maintaining a strong performance in 2024. This suggests that substantial revisions were made to their system. In contrast, teams like MyEachtra and vitrir experienced steep declines in 2024, indicating that their new approaches may not been as effective.
@@
Overall, regular teams exhibited a mix of trends, with some improving significantly over time and others struggling to maintain consistency. It is also important to note that the systems are changing every year to incorporate new approaches and techniques, which can impact performance. The variability in scores across tasks and years highlights the complexity of lifelog retrieval and the need for teams to continuously adapt to remain competitive.
@@ IV-B1 Number of Groundtruth Images
Figure IV-B1 visualises the distribution of the number of groundtruth images associated with each task for KIS and ad-hoc retrieval tasks across the years 2022, 2023, and 2024. Groundtruth images are reference images that are considered correct for a specific query. We do not consider QA tasks in this analysis, as the number of groundtruth images is not applicable to text-based answers.
In KIS tasks, the number of groundtruth images per query remained consistently low in 2022 and 2024, with the median staying close to a small value (under 10). However, in 2023, the distribution shows a significant increase in variability, with some queries having up to around 140–150 groundtruth images. This spike indicates that certain queries in 2023 were associated with longer events from the dataset, making the tasks less about pinpointing a unique image and more about identifying one from a broader set. This may reflect a subtle variation in the nature of the KIS topics that year, with some covering longer time spans than usual, possibly affecting how participants approached the topics. By 2024, the distribution had narrowed again, suggesting a return to queries with more targeted and unique references.
@@
The trends in the number of groundtruth images can help explain some of the system performance results discussed earlier. For KIS tasks, the spike in 2023 groundtruth counts aligns with the higher number of correct submissions and increased precision seen in 2023. Having more reference images per query likely increased the chances of correct matches, making it easier for systems to provide valid answers. This contrasts with 2024, where the reduced number of groundtruth images contributed to a decline in correct submissions and precision. For Ad-hoc tasks, the narrowing of the distribution in 2024 corresponds to the higher performance variability observed among system instances. Queries with fewer groundtruth images require more precise retrieval, which may have exacerbated performance disparities between well-optimised and poorly configured systems. Additionally, if the remaining queries were inherently more difficult or ambiguous, this could explain the increased difficulty in achieving high scores despite improved system tuning.
@@ IV-B2 Groundtruth Coverage and System Performance Across Ad-hoc Tasks
The results show some negative correlation between the Gini coefficient and recall ($r=-0.15$, $p=0.13$) and score ($r=-0.24,p=0.05$), suggesting that unevenly distributed groundtruth images may hinder retrieval performance, even though the effect is relatively weak. This pattern aligns with the scatter plot observations, where tasks with more evenly spread reference images generally performed better than those with clustered timestamps. The correlation between the Gini coefficient and precision was near zero ($r=-0.03$, $p=0.24$), indicating that timestamp distribution appears to affect recall and coverage more than precision.
The number of groundtruth images per task showed a weak negative correlation with recall ($r=-0.17$, $p=0.097$) and score ($r=-0.078$, $p=0.20$), suggesting that retrieval effectiveness may depend more on system design and query formulation than on the sheer availability of reference images.
@@
These findings emphasize the need for future lifelog retrieval systems to adapt their strategies based on both the quantity and temporal distribution of reference images, ensuring more robust and effective retrieval outcomes.

### v2

@@ III Summary of Participating System Approaches
Lifelog retrieval systems must interpret user inputs across multiple modalities, such as free-text queries, visual examples, and contextual filters (e.g., time, location, or activity). The strategies for interpreting these inputs vary by system. Because integration details are team-specific and often unpublished, our discussion focuses on observed usage patterns and system-level effects rather than implementation comparisons.
The strategies used vary significantly across systems, ranging from rule-based filtering to embedding-based semantic search and LLM-driven query interpretation. Tran et al. [78] present a general pipeline for LSC systems, illustrating how user inputs are processed, matched against indexed data, and refined through user interaction. This model reflects the common architecture seen in many LSC submissions and helps contextualise the multimodal input-handling approaches used in the challenge.
The early years of LSC, from 2018 to 2021, saw a gradual progress in lifelog retrieval methods. Developments from this period are detailed in [24] and [81], which documented the advancements made in the initial and later years of this phase, respectively.
@@
This section reviews the major technical advances made in LSC’22, LSC’23, and LSC’24 and provides a structured analysis. Tables III and IV summarise the systems across these editions, highlighting key improvements in UI and UX, search and retrieval techniques, and advanced features.
@@ III-A1 LSC’22: The Rise of CLIP-Based Embedding Models
LSC’22 marked a significant shift toward embedding-based retrieval, with many teams integrating contrastive language–image pretraining (CLIP) models [60] for text-to-image search, which are language-image transformers that are trained in a contrastive way to maximise text-image similarity. Systems such as MyScéal [83] and LifeSeeker 4.0 [53] adopted CLIP to improve semantic understanding, enabling more abstract and flexible queries compared to traditional keyword-based retrieval. Memento 2.0 [1] introduced a weighted ensemble approach for CLIP integration, optimising retrieval efficiency.
User experience improvements were also a focus. MyScéal introduced a more accessible search interface, while LifeSeeker 4.0 used event clustering to enhance navigation. vitrivr-VR [72] introduced expanded their Virtual Reality (VR) search environment, offering three immersive interfaces for lifelog exploration. LIFEXPLORE [44] improved semantic geo-location enrichment and search result ordering to enhance retrieval accuracy. Alternative interaction methods gained traction. Voxento 3.0 [4] implemented voice-based search with speech-to-query mechanisms, while FIRST 3.0 [35] and vitrivr [31] enhanced their retrieval pipelines by incorporating external search engines for unfamiliar concept expansion.
@@ III-B1 Embedding-based Retrieval
Embedding-based retrieval approaches have transformed lifelog search by mapping both text and images into a shared high-dimensional semantic space. Cosine similarity between the search query and the images is directly used to rank the results. Unlike traditional concept-based retrieval, which relies on explicit labels (e.g., ‘beach’ or ‘car’), embeddings capture deeper contextual meanings, enabling the retrieval of relevant data even when exact keywords are missing. For instance, users can issue abstract queries like ‘a relaxing afternoon with friends’ and receive semantically related images and videos.
This technique was mentioned in the previous analysis of LSC’21 [81], and it has become the dominant approach since LSC’22. Text and image embeddings are trained using large datasets of paired image-text data, allowing models such as CLIP [60], BLIP [47], and BLIP-2 [46] to understand complex relationships. CLIP aligns text and image embeddings to maximise their agreement for correct pairs, supporting zero-shot queries. BLIP and BLIP-2 build upon this by improving the visual question-answering and captioning capabilities, making them ideal for multimodal queries. Some optimisation methods, such as KNN search [82] and FAISS [41], can be used to speed up the search as in the case of LIFEXPLORE [69] and Memento [2]. Although there is some effort in fine-tuning the embedding models [77], large-scale pre-trained models are generally more robust and are often used, either directly or in a weighted ensemble [2].
@@
While many systems relied on retrieving images individually, several recent approaches extended these embeddings to incorporate temporal context by operating over structured segments of lifelog data. Rather than treating each image in isolation, systems like MyEachtra [80], Memento 4.0 [3], and MemoriEase 2.0 [86] performed hierarchical or content-based segmentation to construct event-level representations, where visual features from temporally adjacent frames were aggregated. These event embeddings served as the atomic units for retrieval or reasoning, enabling systems to interpret queries involving longer activities or complex temporal dependencies. This shift toward event-centric representations reflects a growing interest in modelling lifelog data in ways that align more closely with human memory and real-world activity flow. Such a trend extends beyond lifelogging and is also evident in the wider multimedia retrieval community, as discussed in [68].
@@ III-B2 Large Language Models, Conversational Search, and Question Answering
The adoption of LLMs in LSC’24 marked a major milestone in lifelog retrieval. LLMs have introduced powerful language understanding capabilities to lifelog retrieval systems. By pretraining on vast corpora of text, LLMs can interpret complex, natural language queries and support dialogue-based search. In lifelog search, this means users can refine their searches through follow-up questions or clarifications, making the retrieval process more interactive and personalised.
@@ III-B3 User Interface and User Experience
  • LSC’23 saw the introduction of dynamic result sequencing and faceted filtering in LIFEXPLORE [69], enhancing search result organisation. MEMORIA [61] and Memento [2] focused on detailed image annotations and user-selectable embedding models, respectively.
  • LSC’24 showcased a range of UI/UX improvements, with several teams participating in the challenge with the focus of UI/UX enhancements for different levels of user expertise, such as Retrospect [74], LifeLens[37], and VitaChronicle [58]. MyEachtraX [84] brought a mobile-friendly design to lifelog search to enhance accessibility. Conversational interfaces, as mentioned earlier, are gaining in popularity, allowing users to interact with lifelog systems through natural language queries. These interfaces further enhance UX by guiding users through the search process step-by-step, allowing for query refinement through dialogue rather than manual re-entry.
Additionally, immersive interfaces gained further traction, with Virtual Reality (VR) and Augmented Reality (AR) lifelog retrieval systems offering novel ways to explore lifelog data. VR allows users to ‘step into’ their lifelogs, navigating time and space through virtual reconstructions. AR overlays lifelog information onto real-world surroundings, enabling contextual re-experiencing of past events. VR systems typically use spatial metaphors such as timelines and maps to organise lifelog entries. Head-gaze tracking and hand gestures often replace traditional input methods, making the interaction more natural. Collaborative VR environments take this further by allowing multiple users to explore lifelog data together. Some notable VR/AR systems showcased at LSC’24 include vitrivr-VR [73], which provided a map-based VR interface for lifelog exploration, and CollaXRSearch [50], which introduced a shared VR workspace for collaborative lifelog analysis. Despite the potential of VR/AR interfaces, they face adoption challenges due to hardware requirements and user comfort. However, as VR/AR technologies mature, they may redefine how users interact with lifelog archives.
@@ IV Comparative Analysis of System Performance through the Years
Since the LSC has been running for several years, it is possible to compare system performance in LSC’24 with previous editions, LSC’22 and LSC’23. The same dataset and tasks were used across all editions, except in 2022, when the QA task required an image as the answer rather than text. Some systems participated across multiple years and experimented with different approaches, giving us insights into the effectiveness of various strategies. This section focuses on trends in system performance across three key metrics: correctness of submissions, time to the first correct submission, and overall scores. We follow a top–down analytical sequence: first, per-metric trends (correctness, timing, scoring); second, query difficulty effects that help explain those trends; and finally, a focused analysis of LSC’24 to illustrate how these factors play out in the most recent edition.
@@ IV-A Performance of Systems
The evaluation metrics used in the LSC are based on the time taken to submit the correct image and the number of correct and incorrect submissions. The number of participating teams increased from 9 in 2022 to 14 in 2023, reaching 35 in 2024. In 2024, some systems had multiple instances (e.g., SYSTEM1, SYSTEM2), which were treated as independent participants. This increased the total number of submissions and led to higher performance variability. We begin with correctness, then examine timing, and finally relate both to aggregate scores across recurring teams.
@@ IV-A1 Correctness of Submissions
For KIS tasks, the mean number of correct submissions increased from 0.78 in 2022 to 0.87 in 2023, indicating improved system strategies. However, it dropped to 0.68 in 2024, while precision decreased to 0.64, likely due to exploratory submissions from system instances with different users. Although some instances were effective, others likely introduced noise, contributing to the lower precision. QA tasks showed consistent correct submissions (0.78 in both 2023 and 2024), but precision improved slightly from 0.50 to 0.52. Ad-hoc tasks exhibited the widest variability, with some teams performing very well and others submitting many incorrect attempts. While the median correct submissions remained stable, the spread widened in 2024, suggesting that not all instances of the same core system performed consistently.
The overall precision for KIS and QA tasks was higher than for ad-hoc tasks, which consistently recorded lower precision due to their open-ended nature. The introduction of multiple system instances in 2024 increased the diversity of approaches, but also amplified the volume of incorrect submissions, impacting precision across tasks. Taken together, these correctness and precision patterns suggest that differences in timing may explain the variability, as slower query reformulation and exploratory submissions likely contribute to it.
@@ IV-A2 Time to First Correct Submission
To contextualise the above correctness patterns, we now analyse time-to-first-correct as a proxy for query formulation and interaction efficiency. Figure IV-A2 shows the distribution of the time taken by teams to make their first correct submission. For KIS and QA tasks, the median times remained relatively stable across years, but the spread increased in 2024 due to system instance variability. This indicates that while some instances responded quickly, others took considerably longer. Outliers in 2024 represent instances that struggled to deliver timely, correct submissions, demonstrating the user variability across instances.
For QA tasks, the spread of times was consistently large across the years, but a slight improvement in the median time in 2024 suggests some gains in system efficiency. The ad-hoc retrieval tasks showed the highest variability in 2024, with some teams submitting correct answers quickly and others taking substantially longer. This wide range of performance reflects the complexity of open-ended queries and the exploratory nature of multiple instances, with different users demonstrating varying levels of success. These timing effects help explain the larger precision differences observed above and prepare for our comparison of scores, which combine correctness and timing in the official evaluation
@@ IV-A3 Scores of Regular Teams by Task Type and Year
In KIS tasks, top-performing teams like MyScéal, LifeInsight, and LifeSeeker maintained consistently high scores across multiple years. However, LifeLens and LifeGraph saw significant score declines in 2024, potentially due to system changes or difficulties adapting to new task types. MEMORIA showed a notable improvement, increasing its score from 10 in 2022 to 90 in 2023, maintaining a strong performance in 2024. This suggests that substantial revisions were made to their system. In contrast, teams like MyEachtra and vitrir experienced steep declines in 2024, indicating that their new approaches may not have been as effective.
@@
Overall, regular teams exhibited a mix of trends, with some improving significantly over time and others struggling to maintain consistency. It is also important to note that the systems are changing every year to incorporate new approaches and techniques, which can impact performance. The variability in scores across tasks and years highlights the complexity of lifelog retrieval and the need for teams to continuously adapt to remain competitive. These score patterns over time motivate a closer look at query difficulty, which may help explain several of the fluctuations noted above.
@@ IV-B Query Difficulty Analysis
We now examine query characteristics to explain when and why systems (and instances) succeed or fail.
@@ IV-B1 Number of Groundtruth Images
Figure IV-B1 visualises the distribution of the number of groundtruth images associated with each task for KIS and ad-hoc retrieval tasks across the years 2022, 2023, and 2024. Groundtruth images are reference images that are considered correct for a specific query. We do not consider QA tasks in this analysis, as the number of groundtruth images does not apply to text-based answers.
In KIS tasks, the number of groundtruth images per query remained consistently low in 2022 and 2024, with the median staying close to a small value (under 10). However, in 2023, the distribution shows a significant increase in variability, with some queries having up to around 140–150 groundtruth images. This spike indicates that certain queries in 2023 were associated with longer events from the dataset, making the tasks less about pinpointing a unique image and more about identifying one from a broader set. This may reflect a subtle variation like the KIS topics that year, with some covering longer time spans than usual, possibly affecting how participants approached the topics. By 2024, the distribution had narrowed again, suggesting a return to queries with more targeted and unique references.
@@
The trends in the number of groundtruth images can help explain some of the system performance results discussed earlier. For KIS tasks, the spike in 2023 groundtruth counts aligns with the higher number of correct submissions and increased precision seen in 2023. Having more reference images per query likely increased the chances of correct matches, making it easier for systems to provide valid answers. This contrasts with 2024, where the reduced number of groundtruth images contributed to a decline in correct submissions and precision. For Ad-hoc tasks, the narrowing of the distribution in 2024 corresponds to the higher performance variability observed among system instances. Queries with fewer groundtruth images require more precise retrieval, which may have exacerbated performance disparities between well-optimised and poorly configured systems. Additionally, if the remaining queries were inherently more difficult or ambiguous, this could explain the increased difficulty in achieving high scores despite improved system tuning. Having established how groundtruth coverage shifted across years, we next consider the temporal distribution of those groundtruths and their link to precision/recall.
@@ IV-B2 Groundtruth Coverage and System Performance Across Ad-hoc Tasks
The results show some negative correlation between the Gini coefficient and recall (r=0.15r=-0.15, $p=0.13$) and score (r=0.24,p=0.05r=-0.24,p=0.05), suggesting that unevenly distributed groundtruth images may hinder retrieval performance, even though the effect is relatively weak. This pattern aligns with the scatter plot observations, where tasks with more evenly spread reference images generally performed better than those with clustered timestamps. The correlation between the Gini coefficient and precision was near zero (r=0.03r=-0.03, $p=0.24$), indicating that timestamp distribution appears to affect recall and coverage more than precision.
The number of groundtruth images per task showed a weak negative correlation with recall (r=0.17r=-0.17, $p=0.097$) and score (r=0.078r=-0.078, $p=0.20$), suggesting that retrieval effectiveness may depend more on system design and query formulation than on the sheer availability of reference images.
@@
These findings emphasise the need for future lifelog retrieval systems to adapt their strategies based on both the quantity and temporal distribution of reference images, ensuring more robust and effective retrieval outcomes. In summary, query difficulty, through both groundtruth size and temporal spread, helps explain the correctness and timing patterns observed earlier. We now look at how these factors appear in LSC’24 specifically.